At what rate can money be invested if $2,320 is equivalent to $2,511.46 paid 186 days later?
Answer by fractalier(6550)

Well if we assume simple interest, then we can use
I = prt
The interest gained was 2511.46 - 2320 = $191.46.
We are looking for r...t = 186/365 years...thus
r = I / pt = (191.46) / (2320)(186/365) = 16.2%